OpenStudy (anonymous):

why is the derivative of \[(Ax+B)e^x\] \[(Ax+A+B)e^x\]? I understand the Ax+A part, but why didn't the B disappear? btw it's an ODE problem. \[y''-4y=xe^x+cos2x\]

OpenStudy (turingtest):

\[y=(Ax+B)e^x\]\[y'=(Ax+B)e^x+Ae^x=(Ax+A+B)e^x\]

OpenStudy (anonymous):

how odd, wolfram alpha won't show me the steps unless I sign up with them.

OpenStudy (anonymous):

why doesn't the B disappear? Isn't it a constant?

OpenStudy (phi):

derivative of B * e^x is B* d/dx (e^x) = B* e^x

OpenStudy (turingtest):

By the product rule the B disappears from one Ax+B term, and in the other term you differentiate e^x \[u=Ax+B\]\[u'=A\]\[v=e^x\]\[v'=e^x\]\[uv'+vu'=(Ax+B)e^x+Ae^x\]factor out \(e^x\) and you get\[(Ax+A+B)e^x\]

OpenStudy (anonymous):

awesome thanks!!!!